# Dressed basis sets for the modeling of exchange interactions in double quantum dots

### v1

@@ III Physics of the exchange interaction
In this section, we discuss the physics of the exchange interaction in Ge/GeSi quantum dots as highlighted by the CI calculations.
@@
The low-energy spectrum of the DQD is plotted in Fig. 7 as a function of the detuning voltage $\delta V_{\mathrm{d}}=2\delta V_{\mathrm{L}}=-2\delta V_{\mathrm{R}}$ with respect to the bias point $V_{\mathrm{L}}=V_{\mathrm{R}}=-40$ mV ($V_{\mathrm{J}}=-15$ mV). This close-up of Fig. 5 highlights the singlets in red and the triplets in blue. The anti-crossings between singlets, but also between triplets are clearly visible on this figure.
We focus on the ground-state singlets to start with. The $S(1,1)$ state anti-crosses the $S(0,2)$ and $S(2,0)$ states at $\delta V_{\mathrm{d}}=\pm\delta V_{\mathrm{d}}^{*}=\pm 8.92$ mV. To convert $\delta V_{\mathrm{d}}$’s into relevant energies, we introduce the detuning lever arm $\alpha_{\mathrm{d}}=\partial\varepsilon/\partial V_{\mathrm{d}}$. From the expectation values of the $D_{\mathrm{L}}$ and $D_{\mathrm{R}}$ operators, we estimate $\alpha_{\mathrm{d}}=0.205$. We can next assess the charging energy $U\approx\alpha_{\mathrm{d}}\delta V_{\mathrm{d}}^{*}=1.83$ meV. This is in good agreement with a direct CI calculation $U=1.97$ meV at $V_{\mathrm{L}}=V_{\mathrm{J}}=0$ (i.e., with the left dot emptied)  [^7] This net charging energy is computed as $U=U_{02}-U_{11}$, where $U_{02}=2.32$ meV is the charging energy in the $(0,2)$ state (beyond the anti-crossing) and $U_{11}=0.346$ meV is the charging in the $(1,1)$ state (at the symmetric operation point). The former is the difference $U_{02}=E[S(0,2)]-E[S^{*}(0,2)]$ between the energies $E[S(0,2)]$ and $E[S^{*}(0,2)]=2E_{0}$ of the interacting and non-interacting singlets, respectively, with $E_{0}$ the single-particle ground-state energy. Likewise, $U_{11}=E[S(1,1)]-E[S^{*}(1,1)]=E[S(1,1)]-2E_{b}$, with $E_{b}$ the bonding ground-state energy at zero detuning.. The anti-crossing gaps at $\delta V_{\mathrm{d}}=\pm\delta V_{\mathrm{d}}^{*}$ are $\Delta_{S}=2|\tau_{S}|=80$$\mu$eV. Therefore, $|\tau_{S}|=40$$\mu$eV is significantly larger than the value $|\tau_{S}^{\mathrm{sp}}|=\sqrt{2}|t|$ expected from the single-particle tunneling $|t|=11.6$$\mu$eV extracted at the $(1,0)/(0,1)$ anti-crossing at zero detuning. The shape of the barrier is indeed different at finite detuning, and, as discussed below, $\tau_{S}$ is altered by Coulomb interactions.
@@
We proceed with a more detailed assessment of Eq. (12). For that purpose, we build the CI basis set from the ground bonding and anti-bonding single-particle states at the symmetric operation point. This ($N=4$, $M=6$) basis set spans the minimal $\{T_{\alpha}(1,1),S(1,1),S(0,2),S(2,0)\}$ subspace acted upon by Hamiltonian (12). We define $J_{0}$ as the exchange energy in this basis set. Once we have identified the singlet/triplet eigenstates of ${\hat{H}}_{\mathrm{CI}}$ and computed $J_{0}$, we diagonalize the detuning operator $D_{\mathrm{d}}=D_{\mathrm{L}}-D_{\mathrm{R}}$ in the singlet subspace in order to split the pure $S(1,1)$ state (with eigenvalue 0) from the pure $S(0,2)$ and $S(2,0)$ states (with eigenvalues ±$\pm d$). We finally rotate ${\hat{H}}_{\mathrm{CI}}$ in the basis set =${\cal B}=\{T_{\alpha}(1,1),S(1,1),S(0,2),S(2,0)\}$ of the pure singlets and triplets and map it to Eq. (12). We extract that way $J_{c}$, $\tau_{S}$, and $U$.
Fig. 6 actually shows that this minimal basis set fails to achieve even qualitative agreement with the converged CI [123, 124]. Its deficiencies nonetheless highlight some key features of the exchange interaction. It turns out that the triplet is the ground-state in the basis set ${\cal B}$, at variance with the theory of Ref. [125] (which only holds, however, in complete basis sets). $J_{c}$ is indeed larger than the “kinetic” exchange $2\tau_{S}^{2}/U$, so that $J_{0}$ is negative. We plot for comparison $J_{c}$ and $-J_{0}$ as a function of $V_{\mathrm{J}}$ in Fig. 8a, as well as the value of $J$ from the CI calculation in the dressed $M^{\prime}=112$ basis set. The net exchange interaction is thus ruled by strong “dynamical” Coulomb correlations outside the quasi-degenerate ground bonding/anti-bonding subspace.
As discussed previously, triplets are favored over singlets by Coulomb interactions owing to the formation of an “exchange hole” in the pair correlation function of the two particles: due to Pauli exclusion principle, two holes in a triplet state can not sit at the same position, while two holes in a singlet can, which strengthens their interaction. The large $J_{c}$ in Fig. 8a originates in the barrier where the two holes leak in the $(1,1)$ states. This is why $J_{c}$ decreases exponentially with $V_{\mathrm{J}}$ as does $t$. The minimal basis set ${\cal B}$ does not, however, allow for correlations between the motions of the two particles (beyond those imposed by anti-symmetry). Yet the two holes practically tend to avoid each other owing to their mutual repulsion. This digs an additional “Coulomb hole” [126] in the pair correlation function of both singlets and triplets, reducing the advantage of the latter. Such dynamical correlations can only be captured in an extended basis set of Slater determinants.
@@
Figure 8: (a) Exchange energies $J$ and $-J_{0}$, and Coulomb exchange correction $J_{c}$ as a function of $V_{\mathrm{J}}$ at the symmetric operation point $\delta V_{\mathrm{d}}=0$. (b) Tunnel couplings $|\tau_{S}|$ and $\sqrt{2}|t|$ as a function of $V_{\mathrm{J}}$ at $\delta V_{\mathrm{d}}=0$.
Moreover, the effective tunneling strength between the two dots is significantly affected by Coulomb correlations. The effect of Coulomb interactions on the coupling between the $(1,1)$ and $(2,0)/(0,2)$ states is already sizeable in the minimal basis set ${\cal B}$. We compare in Fig. 8b the value of $\tau_{S}$ extracted in basis set ${\cal B}$ with the expected $\sqrt{2}t$ from Fig. 3. They differ significantly over the whole $V_{\mathrm{J}}$ range. Indeed, the Coulomb interaction has non zero matrix elements between the $S(1,1)$ and $S(0,2)/S(2,0)$ states, and even between the $S(0,2)$ and $S(2,0)$ states themselves (although they are all much smaller than the charging energy $U$). The Coulomb interaction thus plays a key role in shaping the net interactions between the $(1,1)$ and $(0,2)/(2,0)$ charge sectors.
@@
Another quantitative flaw of Eq. (12) is that it misses the curvature of the triplet states [127, 104]. Indeed, the triplet $T(1,1)$ state anti-crosses $T(2,0)$ and $T(0,2)$ states at higher detuning and thus also bends down, which reduces the net exchange energy but is not accounted for by Eq. (12). This correction is controlled by the gap $\Delta_{ST}$ between the $S(0,2)$ and $T(0,2)$ states (or between the $S(2,0)$ and $T(2,0)$ states), and by the triplet tunneling strength $\tau_{T}$. The Coulomb interaction has naturally a strong impact on the $(0,2)$ states, and leads, in particular, to a renormalization of $\Delta_{ST}$. In a naive, weakly interacting picture, $\Delta_{ST}^{0}=\varepsilon_{1}-\varepsilon_{0}$ is the difference between the first excited ($\varepsilon_{1}$) and ground ($\varepsilon_{0}$) single-particle energy levels of the right quantum dot (as a triplet can not be built from a single level). As discussed above, Coulomb interactions however favor triplets over singlets, which lowers the actual splitting $\Delta_{ST}<\Delta_{ST}^{0}$ [128]. This reduction is significant in the present device. A CI calculation with the left dot empty at $V_{\mathrm{L}}=0$ V and $V_{\mathrm{J}}=0$ V indeed yields $\Delta_{ST}=0.47$ meV while $\Delta_{ST}^{0}=1.32$ meV. This estimate of $\Delta_{ST}$ is roughly consistent with the splitting between the singlet and triplet anti-crossings on Fig. 7, $\Delta V_{\mathrm{d}}=2$ mV or $\Delta\varepsilon=0.41$ meV. The renormalization of $\Delta_{ST}$ is a key ingredient of the physics of the exchange interaction, as a smaller $\Delta_{ST}$ increases the curvature of the triplet states in the $(1,1)$ sector and lowers the exchange energy. We remind that $\Delta_{ST}$ may collapse in more elongated quantum dots due to Wigner localization effects [99, 100, 101, 102, 103, 104, 105, 106, 107, 108]: Coulomb correlations localize the two holes one on each side of the dot, so that the density looks alike in the $S(0,2)$ and $T(0,2)$ states (which have the fingerprints of an effective $(1,1)$ state built in the right dot). $\Delta_{ST}$ then decreases almost exponentially with the aspect ratio of the dot. The effects of Coulomb correlations increase with the effective mass, which makes germanium more resilient to Wigner localization than, e.g., silicon. The gap $2\tau_{T}$ is also visibly larger at the $T(1,1)/T(0,2)$ than at the $S(1,1)/S(0,2)$ anti-crossing on Fig. 7, which further enhances triplet curvature. This results from the different symmetry and extension of the states, and from an overall reduction of the effective barrier height when increasing detuning.
@@ IV Effects of the magnetic field
In the absence of SOC, the magnetic field simply splits the threefold degenerate triplet into $T_{0}$, $T_{+}$ and $T_{-}$ states. Vertical confinement and SOC however split the $p$-like Bloch functions into HH, LH and split-off subbands with different angular momenta and $g$-factors. The admixture of HH and LH components by lateral confinement and strains modulates the net $g$-factors that can be different in the two dots [46, 48, 50, 53]. Additionally, SOC gives rise to Rashba- and Dresselhaus-like interactions that couple the pseudo-spin and orbital motion (envelopes) of the hole [42, 45, 47, 50, 51]. This mixes the split singlet and triplet states and gives rise to a complex pattern of anti-crossings as a function of magnetic field strength and orientation [59, 60, 62, 61].
The effects of SOC on coupled quantum dots can be qualitatively understood in the minimal $\{S(2,0),S(0,2),S(1,1),T_{0}(1,1),T_{+}(1,1),T_{-}(1,1)\}$ subspace. For that purpose, we introduce the ground-state orbitals $\lvert L\uparrow\rangle$ and $\lvert L\downarrow\rangle$ of the left dot, as well as the ground-state orbitals $\lvert R\uparrow\rangle$ and $\lvert R\downarrow\rangle$ of the right dot. The pseudo-spin index $\uparrow$/$\downarrow$ / ↓ labels the Kramers-degenerate left and right orbitals at $B=0$, and may be chosen differently in the two dots. We next assume the following one-particle Hamiltonian in the $\{\lvert L\uparrow\rangle,\lvert L\downarrow\rangle,\lvert R\uparrow\rangle,\lvert R\downarrow\rangle\}$ basis set [79]
@@
[…]
@@
We emphasize that $t_{\mathrm{sc}}$ and $t_{\mathrm{sf}}$ depend on the orientation of the magnetic field $\bm{b}$ (although $t=\sqrt{|t_{\mathrm{sc}}|^{2}+|t_{\mathrm{sf}}|^{2}}$ does not), and that $t_{\mathrm{sc}}$ can always be made real by an appropriate choice of phases for the $L$ and $R$ orbitals. In fact, $|t_{\mathrm{sc}}|$ and $|t_{\mathrm{sf}}|$ have very simple expressions as a function of the $g$-matrices $\tilde{g}_{L}$ and $\tilde{g}_{R}$ in a spin-orbit frame (see also note [122]):
@@
[…]
@@
[…]
@@
As an illustration, we plot in Figs. 10 and 11 the maps of $J_{\parallel}$ and $\Delta_{ST_{-}}$ as a function of the orientation of the magnetic field in two relevant cases: the same quasi-circular dots as in the previous figures (B gates grounded), and slightly squeezed dots ($V_{\mathrm{B}_{1}}=V_{\mathrm{B}_{5}}=50$ mV and $V_{\mathrm{B}_{2}}=V_{\mathrm{B}_{6}}=-15$ mV)  [^9] For the squeezed dots, $\delta V_{\mathrm{d}}=\delta V_{\mathrm{L}}-\delta V_{\mathrm{R}}$ with $\delta V_{\mathrm{R}}=-1.035\delta V_{\mathrm{L}}$. . The inhomogeneous strains brought by the differential thermal contraction of the metals and insulators are taken into account in these calculations. The maps are computed from the dependence of the ground-state singlet and triplet energies on the magnetic field amplitude, illustrated in Fig. 9 at specific magnetic field orientations. The field $B_{ST_{-}}$ and gap $\Delta_{ST_{-}}$ of the $ST_{-}$ anti-crossing are found by dichotomy, then the anisotropic exchange energy is estimated as $J_{\parallel}=E_{T^{+}}+E_{T^{-}}-E_{\Uparrow\Downarrow}-E_{\Downarrow\Uparrow}$ at magnetic field $B=4B_{ST_{-}}$. The use of the dressed basis set enables efficient extractions of these quantities and the production of very detailed maps.
Table 1: Fitted values for the tunnel couplings ($\mu$eV) and $g$-matrices of the circular and squeezed dots.
@@
Circular dots $-8.27$ $-0.055$ $0.173$ $0.147$ $-0.104$ $-0.115$ $14.23$ $14.19$ $-0.0143$ $0.0123$
Squeezed dots $-11.3$ $-0.084$ $0.229$ $0.120$ $-0.020$ $-0.141$ $13.32$ $12.82$ $-0.0141$ $0.0177$
@@
The tunnel coupling $t_{2}$ is more than two orders of magnitude smaller than $t_{1}$. This suggests that the Rashba SOC is weak in the present devices. Its strength can alternatively be characterized by the spin-orbit length $\ell_{\mathrm{so}}$ such that the pseudo-spin of the hole rotates by an angle $\alpha=2d/\ell_{\mathrm{so}}$ when tunneling over a distance $d$ [136, 137, 132]. This spin-orbit length can be estimated as $\ell_{\mathrm{so}}\approx a_{\mathrm{2D}}|t_{1}/t_{2}|\approx 25\,\mu$m, which is much longer than the inter-dot separation $a_{\mathrm{2D}}=180$ nm. As a matter of fact, Rashba interactions arise from the breaking of the inversion symmetry by the structure and applied electric field and by the inhomogeneous shear strains [24, 42, 50]. However, the cubic Rashba interaction is small at low vertical electric fields in such devices [51]. Moreover, the linear Rashba interaction arising from inhomogeneous shear strains is almost zero on average on the way from one dot to the other as the gradient of $\varepsilon_{xz}$ changes sign between the dots [50].
Therefore, the main features of Figs. 10 and 11 are essentially preserved if $t_{2}$ is neglected, in which case Eq. (20) holds with the $g$-matrices of Table 1. $t_{\mathrm{sf}}$ thus primarily results from the mismatch between the principal $g$-factors and axes of the two dots. It is almost zero when the magnetic field is along the principal $\approx\bm{x}$, $\bm{y}$ or $\approx\bm{z}$ axes (because the pseudo-spin precession axes $\bm{\omega}_{L}$ and $\bm{\omega}_{R}$ remain aligned despite the different principal $g$-factors). Moreover, $\bm{\omega}_{L}$ and $\bm{\omega}_{R}$ get locked onto the $\bm{z}$ axis once the magnetic field goes out-of-plane owing to the large $g_{zz}\gg g_{xx},\,|g_{yy}|$. Consequently, $t_{\mathrm{sf}}$, $\Delta_{ST_{-}}$ and the deviations $J-J_{\parallel}$ are significant only near the equatorial plane. Leaving out $g_{zx}$, $|t_{\mathrm{sf}}|$ is actually maximal when $\theta\approx 90^{\circ}$ and  [^10] These expressions apply for positive $g_{xx}$’s and negative $g_{yy}$’s.
@@
Using these expressions, we find $|t_{\mathrm{sf}}|_{\mathrm{max}}=0.53\,\mu$eV at $\varphi=\pm(90\pm 34.49)^{\circ}$ in the quasi-circular dots and $|t_{\mathrm{sf}}|_{\mathrm{max}}=5.63\,\mu$eV at $\varphi=\pm(90\pm 17.62)^{\circ}$ in the squeezed dots, in fair agreement with the estimates drawn from Eq. (23) on Figs. 10 and 11. The effects of SOC are thus stronger in the squeezed dots where the mismatch between the principal $g$-factors is larger.

### v2

@@ III Physics of the exchange interaction
In this section, we discuss the physics of the exchange interaction in Ge/GeSi quantum dots. We first introduce the minimal CI Hamiltonian for exchange (the Hund-Mulliken model). We then discuss the shortcomings of this minimal model by comparisons with CI calculations for large $M$’s. We show, in particular, that it misses relevant excitations as well as dynamical correlations. They are actually captured by the dressed basis set at a reasonable cost.
@@
The low-energy spectrum of the DQD is plotted in Fig. 7 as a function of the detuning voltage $\delta V_{\mathrm{d}}=2\delta V_{\mathrm{L}}=-2\delta V_{\mathrm{R}}$ with respect to the bias point $V_{\mathrm{L}}=V_{\mathrm{R}}=-40$ mV ($V_{\mathrm{J}}=-15$ mV). This close-up of Fig. 5 highlights the singlets in red and the triplets in blue. The anti-crossings between the singlet states are clearly visible on this figure; however, Eq. (12) obviously misses the nearby anti-crossings between the $T(1,1)$ and $T(0,2)/T(2,0)$ states, whose effects on the exchange energy will be addressed later.
[…]
@@
We proceed with a more detailed assessment of Eq. (12). For that purpose, we construct the CI basis set from the ground bonding and anti-bonding single-particle states of the LK Hamiltonian at the symmetric operation point. The $M=6$ Slater determinants built from these $N=4$ bonding and anti-bonding states are actually linear combinations of the $\{T_{\alpha}(1,1),S(1,1),S(0,2),S(2,0)\}$ states acted upon by the Hund-Mulliken Hamiltonian (they span the same subspace). We define $J_{0}$ as the exchange energy in this CI basis set. Once we have identified the singlet/triplet eigenstates of ${\hat{H}}_{\mathrm{CI}}$ and computed $J_{0}$, we diagonalize the detuning operator $D_{\mathrm{d}}=D_{\mathrm{L}}-D_{\mathrm{R}}$ in the singlet subspace in order to split the pure $S(1,1)$ state (with eigenvalue 0) from the pure $S(0,2)$ and $S(2,0)$ states (with eigenvalues $\pm d$). We finally rotate ${\hat{H}}_{\mathrm{CI}}$ in the basis set ${\cal B}=\{T_{\alpha}(1,1),S(1,1),S(0,2),S(2,0)\}$ of the pure singlets and triplets and map it to Eq. (12). We extract that way $J_{c}$, $\tau_{S}$, and $U$.
[…]
@@
Figure 8: (a) Exchange energies $J$ and $J_{0}$, and Coulomb exchange correction $J_{c}$ as a function of $V_{\mathrm{J}}$ at the symmetric operation point $\delta V_{\mathrm{d}}=0$. (b) Tunnel couplings $|\tau_{S}|$ and $\sqrt{2}|t|$ as a function of $V_{\mathrm{J}}$ at $\delta V_{\mathrm{d}}=0$.
[…]
@@
Another quantitative flaw of Eq. (12) is that it misses the curvature of the triplet states [128, 104]. Indeed, the triplet $T(1,1)$ state anti-crosses $T(2,0)$ and $T(0,2)$ states at higher detuning and thus also bends down, which reduces the net exchange energy but is not accounted for by Eq. (12). This correction is controlled by the gap $\Delta_{ST}$ between the $S(0,2)$ and $T(0,2)$ states (or between the $S(2,0)$ and $T(2,0)$ states), and by the triplet tunneling strength $\tau_{T}$. The Coulomb interaction has naturally a strong impact on the $(0,2)$ states, and leads, in particular, to a renormalization of $\Delta_{ST}$. In a naive, weakly interacting picture, $\Delta_{ST}^{0}=\varepsilon_{1}-\varepsilon_{0}$ is the difference between the first excited ($\varepsilon_{1}$) and ground ($\varepsilon_{0}$) single-particle energy levels of the right quantum dot (as a triplet can not be built from a single level). As discussed above, Coulomb interactions however favor triplets over singlets, which lowers the actual splitting $\Delta_{ST}<\Delta_{ST}^{0}$ [129]. This reduction is significant in the present device. A CI calculation with the left dot empty at $V_{\mathrm{L}}=0$ V and $V_{\mathrm{J}}=0$ V indeed yields $\Delta_{ST}=0.47$ meV while $\Delta_{ST}^{0}=1.32$ meV. This estimate of $\Delta_{ST}$ is roughly consistent with the splitting between the singlet and triplet anti-crossings on Fig. 7, $\Delta V_{\mathrm{d}}=2$ mV or $\Delta\varepsilon=0.41$ meV. The renormalization of $\Delta_{ST}$ is a key ingredient of the physics of the exchange interaction, as a smaller $\Delta_{ST}$ increases the curvature of the triplet states in the $(1,1)$ sector and lowers the exchange energy. We remind that $\Delta_{ST}$ may collapse in elongated quantum dots due to Wigner localization effects [99, 100, 101, 102, 103, 104, 105, 106, 107, 108]: Coulomb correlations localize the two holes one on each side of the dot, so that the density looks alike in the $S(0,2)$ and $T(0,2)$ states (which have the fingerprints of an effective $(1,1)$ state built in the right dot). $\Delta_{ST}$ then decreases almost exponentially with the aspect ratio of the dot. Nevertheless, the effects of Coulomb correlations scale with the effective mass, which makes germanium more resilient to Wigner localization than, e.g., silicon. The gap $2\tau_{T}$ is also visibly larger at the $T(1,1)/T(0,2)$ than at the $S(1,1)/S(0,2)$ anti-crossing on Fig. 7, which further enhances triplet curvature. This results from the different symmetry and extension of the states, and from an overall reduction of the effective barrier height when increasing detuning.
@@ IV Effects of the magnetic field
In the absence of SOC, the magnetic field simply splits the threefold degenerate triplet into $_{0}, T+T_{+}$ and $T_{-}$ states. Vertical confinement and SOC however split the $p$-like Bloch functions into HH, LH and split-off subbands with different angular momenta and $g$-factors. The admixture of HH and LH components by lateral confinement and strains modulates the net $g$-factors that can be different in the two dots [46, 48, 50, 53]. Additionally, SOC gives rise to Rashba- and Dresselhaus-like interactions that couple the pseudo-spin and orbital motion (envelopes) of the hole [42, 45, 47, 50, 51]. This mixes the split singlet and triplet states and gives rise to a complex pattern of anti-crossings as a function of magnetic field strength and orientation [59, 60, 62, 61].
The effects of SOC on coupled quantum dots can be qualitatively understood in the minimal $\{S(2,0),S(0,2),S(1,1),T_{0}(1,1),T_{+}(1,1),T_{-}(1,1)\}$ subspace. For that purpose, we introduce the ground-state orbitals $\lvert L\uparrow\rangle$ and $\lvert L\downarrow\rangle$ of the left dot, as well as the ground-state orbitals $\lvert R\uparrow\rangle$ and $\lvert R\downarrow\rangle$ of the right dot. The pseudo-spin index $\uparrow/\downarrow$ labels the Kramers-degenerate left and right orbitals at $B=0$, and may be chosen differently in the two dots. We next assume the following one-particle Hamiltonian in the $\{\lvert L\uparrow\rangle,\lvert L\downarrow\rangle,\lvert R\uparrow\rangle,\lvert R\downarrow\rangle\}$ basis set [79]
@@
is, nonetheless, independent on the choice of pseudo-spins. Practically, the pseudo-spins can always be chosen so that $T=tI_{2}$ is diagonal [74]. In such a “spin-orbit frame”  [^8] The spin-orbit frame is not uniquely defined either, as a same unitary transform on both dots leaves $T$ invariant. A particular spin-orbit frame can always be built starting from given $g_{L}$, $g_{R}$ matrices and tunneling block $T$ parametrized as $T=t(\cos\theta_{\mathrm{so}}I_{2}-i\sin\theta_{\mathrm{so}}\bm{n}_{\mathrm{so}}\cdot\bm{\sigma})$, with $\bm{n}_{\mathrm{so}}$ and $\theta_{\mathrm{so}}$ the so-called spin-orbit vector and spin-orbit angle. Indeed, the transformations $U_{L}=\exp(-i\theta_{\mathrm{so}}\bm{n}_{\mathrm{so}}\cdot\bm{\sigma}/2)$ in the $\{\lvert L\uparrow\rangle,\lvert L\downarrow\rangle\}$ subspace and $U_{R}=\exp(i\theta_{\mathrm{so}}\bm{n}_{\mathrm{so}}\cdot\bm{\sigma}/2)$ in the $\{\lvert R\uparrow\rangle,\lvert R\downarrow\rangle\}$ subspace diagonalize the $T$ block $\tilde{T}=U_{L}^{\dagger}TU_{R}=tI_{2}$. The $g$-matrices then become $\tilde{g}_{L}=R(\bm{n}_{\mathrm{so}},-\theta_{\mathrm{so}})g_{L}$ and $\tilde{g}_{R}=R(\bm{n}_{\mathrm{so}},\theta_{\mathrm{so}})g_{R}$ with $R(\bm{u},\theta)$ the matrix of the real-space rotation of angle $\theta$ around the unit vector $\bm{u}$ [74, 61]., the effects of SOC are fully embedded in the $g$-matrices $\tilde{g}_{L}$ and $\tilde{g}_{R}$ (whose analysis may, however, be intricate). Alternatively, $\lvert\uparrow\rangle$ and $\lvert\downarrow\rangle$ may be chosen, respectively, as the states with largest $\lvert+\tfrac{3}{2}\rangle$ and $\lvert-\tfrac{3}{2}\rangle$ components in each dot. In that case, the effects of SOC are distributed between the $g$-matrices and $T$ block, which are, however, more easily analyzed against perturbation theories of pure HH states [48, 50]. The observables are independent on this basis choice, including the effective $g$-factors of the left dot, $g_{L}^{*}=|g_{L}\bm{b}|$, and right dot, $g_{R}^{*}=|g_{R}\bm{b}|$, with $\bm{b}=\bm{B}/B$ the unit vector along the magnetic field.
@@
We emphasize that $t_{\mathrm{sc}}$ and $t_{\mathrm{sf}}$ depend on the orientation $\bm{b}$ of the magnetic field (although $t=\sqrt{|t_{\mathrm{sc}}|^{2}+|t_{\mathrm{sf}}|^{2}}$ does not), and that $t_{\mathrm{sc}}$ can always be made real by an appropriate choice of phases for the $L$ and $R$ orbitals. In fact, $|t_{\mathrm{sc}}|$ and $|t_{\mathrm{sf}}|$ have very simple expressions as a function of the $g$-matrices $\tilde{g}_{L}$ and $\tilde{g}_{R}$ in a spin-orbit frame (see also note [132]):
@@
where $\delta g^{*}=(g_{\mathrm{L}}^{*}-g_{\mathrm{R}}^{*})/2$ is the half difference and $\bar{g}^{*}=(g_{\mathrm{L}}^{*}+g_{\mathrm{R}}^{*})/2$ the average of the $g$-factors of the dots. When $B\to 0$, this Hamiltonian is the same (up to a unitary transform) as Eq. (12) with $t=\sqrt{t_{\mathrm{sc}}^{2}+|t_{\mathrm{sf}}|^{2}}$ and $J_{c}$ neglected. The magnetic field splits the $T_{+}(1,1)$ and $T_{-}(1,1)$ states by $\Delta E=2\mu_{B}B\bar{g}^{*}$, and mixes the $S(1,1)$ and $T_{0}(1,1)$ states with strength $\Delta E_{Z}=2\mu_{B}B\delta g^{*}$. This gives rise to a competition between the Zeeman and exchange interactions. At very low magnetic fields, the ground-state is the mixed $S(1,1)/S(0,2)/S(2,0)$ singlet split from the $T_{0}(1,1)$ triplet by the exchange energy $J$. The magnetic field tends to break this singlet into anti-symmetric $\lvert\Uparrow\Downarrow\rangle$ and $\lvert\Downarrow\Uparrow\rangle$ states, which become the approximate eigenstates at large $\Delta E_{Z}\gg J$. In the regime $0\ll\varepsilon\ll U$ [where $S(1,1)$ weakly mixes with $S(0,2)$] and at high magnetic field $\Delta E_{Z}\gg J$, the eigenenergies are actually
@@
is the net exchange interaction resulting from the $S(1,1)/S(0,2)$ mixing by $t_{\mathrm{sc}}$ and the $T_{\pm}(1,1)/S(0,2)$ mixing by $t_{\mathrm{sf}}$ (with $\delta\varepsilon=U-\varepsilon$). Note that the exchange splitting $J_{\parallel}$ at high magnetic field is generally different from $J$ (unless $t_{\mathrm{sf}}=0$) and dependent (like $t_{\mathrm{sc}}$ and $t_{\mathrm{sf}}$) on the orientation of the magnetic field [77, 74]. The crossover between the low and high field regimes is well visible in Fig. 9b, which displays the singlet and triplet energies as a function of the magnetic field amplitude in a squeezed double dot with different $g$-factors (see details below). The exchange shift $J_{\parallel}/2$ of the $\lvert\Uparrow\Downarrow\rangle$ and $\lvert\Downarrow\Uparrow\rangle$ states can be exploited to perform, e.g., conditional rotations [19].
@@
As an illustration, we plot in Figs. 10 and 11 the maps of $J_{\parallel}$ and $\Delta_{ST_{-}}$ as a function of the orientation of the magnetic field in two relevant cases: the same quasi-circular dots as in the previous figures (B gates grounded), and slightly squeezed dots ($V_{\mathrm{B}_{1}}=V_{\mathrm{B}_{5}}=50$ mV and $V_{\mathrm{B}_{2}}=V_{\mathrm{B}_{6}}=-15$ mV)  [^9] For the squeezed dots, $\delta V_{\mathrm{d}}=\delta V_{\mathrm{L}}-\delta V_{\mathrm{R}}$ with $\delta V_{\mathrm{R}}=-1.035\delta V_{\mathrm{L}}$.. We include here the inhomogeneous strains imposed by the differential thermal contraction of metals and insulators, in order to unveil their effects on the exchange interaction and $ST_{-}$ splitting. The maps are computed from the dependence of the ground-state singlet and triplet energies on the magnetic field amplitude, illustrated in Fig. 9 at specific magnetic field orientations. The field $B_{ST_{-}}$ and gap $\Delta_{ST_{-}}$ of the $ST_{-}$ anti-crossing are found by dichotomy, then the anisotropic exchange energy is estimated as $J_{\parallel}=E_{T^{+}}+E_{T^{-}}-E_{\Uparrow\Downarrow}-E_{\Downarrow\Uparrow}$ at magnetic field $B=4B_{ST_{-}}$. The use of the dressed basis set enables efficient extractions of these quantities and the production of very detailed maps.
[…]
@@
Circular dots 8.27-8.27 0.055-0.055 $0.173$ $0.147$ 0.104-0.104 0.115-0.115 $14.23$ $14.19$ 0.0143-0.0143 $0.0123$
Squeezed dots 11.3-11.3 0.084-0.084 $0.229$ $0.120$ 0.020-0.020 0.141-0.141 $13.32$ $12.82$ 0.0141-0.0141 $0.0177$
@@
[…]
Therefore, the main features of Figs. 10 and 11 are essentially preserved if $t_{2}$ is neglected, in which case Eq. (20) holds with the $g$-matrices of Table 1. $t_{\mathrm{sf}}$ thus primarily results from the mismatch between the principal $g$-factors and (to a lesser extent) magnetic axes of the two dots. It is almost zero when the magnetic field is along the principal $\approx\bm{x}$, $\bm{y}$ or $\approx\bm{z}$ axes (because the pseudo-spin precession axes $\bm{\omega}_{L}$ and $\bm{\omega}_{R}$ remain aligned despite the different principal $g$-factors). Moreover, $\bm{\omega}_{L}$ and $\bm{\omega}_{R}$ get locked onto the $\bm{z}$ axis once the magnetic field goes out-of-plane owing to the large $g_{zz}\gg g_{xx},\,|g_{yy}|$. Consequently, $t_{\mathrm{sf}}$, $\Delta_{ST_{-}}$ and the deviations $J-J_{\parallel}$ are significant only near the equatorial plane. Leaving out $g_{zx}$, $|t_{\mathrm{sf}}|$ is actually maximal when $\theta\approx 90^{\circ}$ and  [^10] These expressions apply for positive $g_{xx}$’s and negative $g_{yy}$’s.
@@
[…]